OpenStudy (anonymous):

finding the slope of an equation of a line. -1=-3x + y

OpenStudy (nirmalnema):

slope=-a/b =-(-3)/1 =3

OpenStudy (irishboy123):

try adding 3x to each side of the equation and see what happens!

OpenStudy (anonymous):

turn it into y=mx+b where m is the slope, y=3x-1 ....^ slope=3

OpenStudy (anonymous):

thank you guys.!

OpenStudy (anonymous):

no prob